OpenStudy (anonymous):

5/2a-4/is greater than or equal to 15 solve

OpenStudy (ghazi):

\[\frac{ 5 }{ 2 }a-4 \le 15\] is this your question?

OpenStudy (anonymous):

no it's: \[5|2a-4| \ge15\]

OpenStudy (ghazi):

I2a-4I<= 3 -3<2a-4<3 add 4 everywhere lol -3+4<2a<3+4 1<2a<7 divide everywhere by 2 0.5< a< 3.5

OpenStudy (anonymous):

what happens to the 5?

OpenStudy (ghazi):

we divide both the sides by 5 initially

OpenStudy (anonymous):

im confused :p how did you do it?

OpenStudy (ghazi):

5|2a−4|≥15 just divide both the sides by 5 5 I2a-4I /5 => 15/5 I2a-4I=> 3

OpenStudy (anonymous):

then you add 4?

OpenStudy (ghazi):

correct :) so in the middle -4+4=0

OpenStudy (anonymous):

so it becomes 2a>7 ?

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

no you are taking mods off so let say we have IxI>4 then it will break like -4<x<4

OpenStudy (anonymous):

so what happens next then?

OpenStudy (ghazi):

then you have the whole solution right up :D

OpenStudy (anonymous):

whatt? dont you have to keep solve for a? :P

OpenStudy (ghazi):

i have solved for a a>0.5 , a<3.5

OpenStudy (anonymous):

thank you! <3